the dimension of a closed box is 30cm by 50cm by 10cm high.if the thikness of six panels is 1cm find its volume
Answer by checkley77(12569)

BASE=50*30=1.500 CM^2
IF THE 10 CM HIGH SIDES ARE OUTSIDE THE 50 BY 30 BASE THEN THE HEIGHT IS 8 BECAUSE YOU HAVE TO SUBTRACT 2 CM TO ACCOUNT FOR THE THICKNESS OF THE BASE AND LID:
50 by 30
|---------------|
|***********|
|***********| 8
|***********|
|---------------|
1,500*8=12,000 CM^3 IS THE VOLUME.
IF THE 10 CM HIGH SIDES SIT ON THE BASE THEN THE BASE IS 48 BY 28 BECAUSE YOU HAVE TO SUBTRACT 2 CM. FROM THE BASE AND LID MEASUREMENTS TO ACCOUNT FOR THE THICKNESS OF THE SIDES:
50 by 30
_________
|********|
|********|
|********| 10
|********|
_________
SORRY THAT THE SKETCHES AREN'T BETTER BUT THAT'S THE BEST I CAN DO WITH THE KEYBOARD SYMBOLS.
(50-2)(30-2)*10=48*28*10=13,440 CM^3 IS THE VOLUME.